Find the component of vector A=ax i^ +ay j^ + az k^ along the direction of i^-j^

Grade:11

1 Answers

Khimraj
5 years ago
 magnitude of component of A along \hat{i} - \hat{j} = A.(\hat{i} - \hat{j} ) = ax -ay
direction of component = [(\hat{i} - \hat{j} )/\sqrt{2}]
Then component = (ax\hat{i} - ay\hat{j} )/\sqrt{2}
Hope it clears.